epiclearner:

Chloe is playing the flute some distance away from a crowd. If the atmospheric temperature is 15°C, what is the speed of the sound traveling from the flute?

Shadow:

Do you know of a formula for solving this?

Shadow:

Just a moment

Shadow:

Do you know how to convert from Celsius to Kelvin?

epiclearner:

no

Shadow:

I can give you the formula

Shadow:

\[T _{k} = T _{c} + 273.15\] Where tc = the temperature in celsius and tk = the temperature in kelvin

Shadow:

So what would you do?

epiclearner:

plug in 15c into Tk?

Shadow:

No, 15 is the temperature in Celsius, so we would plug it into Tc

Shadow:

Then solve for Tk

epiclearner:

thats what i meant

Shadow:

What do you get for Tk?

epiclearner:

I add 15 to 273.15

Shadow:

288.15 right?

epiclearner:

correct

Shadow:

So the reason we converted from Celsius, to Kelvin, is because the formula for calculating the speed of sound only works with Kelvins

Shadow:

\[V _{sound} = 645.855 \times (\frac{ T _{k} }{ 273.15 })^{0.5}\] Where V sound = the speed of sound, and Tk as we know, is the temperature in Kelvins

Shadow:

You can also look at the formula this way: \[v _{sound} = 645.855 \times \sqrt (\frac{ T _{k} }{ 273.15 })\]

Shadow:

Let me know if you need help

epiclearner:

i need hep

Shadow:

What part are you confused with?

epiclearner:

I am just confused on how they turned it into the hertz equation

Shadow:

What do you mean/

Shadow:

?

epiclearner:

because the answer choices are like 5.52 x 10 hertz

Shadow:

Interesting

Shadow:

Can you provide me with all of the answer choices? Perhaps a screenshot?

epiclearner:

IM on my laptop

Shadow:

There is a screenshot tool by the name of LightShot. It's very easy and simple, works for Mac/Windows.

Shadow:

Let me know if you can get it. If not we can figure it out somehow.

epiclearner:

i don't know im sorry

Shadow:

This is the LightShot tool in action. http://prntscr.com/iysupu If you followed the equation, you would have gotten 340.2 m/s

Shadow:

I looked up examples of this being solved and the speed of sound was always given in m/s.

Shadow:

hertz sounds like frequency, which is a bit different